Victoria Korchnoi is thinking of importing caviar to sell to restaurants and specialty stores. She estimates that this venture will require an initial outlay of $600,000 to buy a refrigerated storage unit, which can be depreciated (straight-line) to a salvage value of $100,000 in 15 years. In addition, Ms. Korchnoi estimates that she will need $70,000 in working capital during the 15 years of the project. Annual sales are estimated to be $180,000 and annual expenses $45,000. At the end of the tenth there will be an expense of $100,000 to properly maintain the storage unit. Ms. Korchnoi estimates that the marginal tax rate will be 30% during the lifetime of the project.

(a) What is the initial outlay associated with starting the business?

(b) What is the annual cash flow from operations?

(c) What is the payback period for this project?

(d) If Ms. Korchnoi requires a 16% RRR to make this investment, should she make the investment.

(e) What is the project's IRR?
Answer #1

a]

Initial outlay = cost of refrigerated storage unit + investment in working capital

Initial outlay = $600,000 + $70,000 =  $670,000

b]

Operating cash flow (OCF) each year = income after tax + depreciation

depreciation each year = (cost of equipment - salvage value) / life of project

depreciation each year = ($600,000 - $100,000) / 15 = $33,333

In year 15, the entire working capital investment is recovered.

In year 15, the equipment is sold for its salvage value of $100,000. There is no tax effect as the book value equals the salvage value.

Annual cash flows are shown in the table below

c]

Payback period is the time taken for the cumulative cash flows to equal zero

Payback period = 6 + (cash flow required in year 7 for cumulative cash flows to equal zero / year 7 cash flow) = 6 + ( $43,000 / $104,500) = 6.41 years

d]

NPV is calculated using NPV function in Excel

NPV is -$84,885

She should not make the investment as the NPV is negative

e]

IRR is calculated using IRR function in Excel

IRR is 13.33%